OpenStudy (anonymous):

James is going to borrow 14000 from his credit union to buy a used car. The apr is 7.0% and the length of the loan is 4 years how much will jamal pay in intrest?

OpenStudy (lilg132):

14000/100 = 140 140 * 7 = 980 interest per year 980 * 4 = $3290 jamal will have to pay in interest remember to click good answer